In a hospital’s shipment of 2,880 insulin syringes, 14 were unusable due to defects. (a) At α = .05,
is this sufficient evidence to reject future shipments from this supplier if the hospital’s quality
standard requires 99.7 percent of the syringes to be acceptable? State the hypotheses and decision
rule.(d) Find the p-value.
I have tried to work on this problem and this is what I got as the answer -16.98797045
please help me.

Answer by stanbon(75887)

In a hospital’s shipment of 2,880 insulin syringes, 14 were unusable due to defects.
p-hat = 2866/2880 = 0.995
--------------------------
(a) At α = .05, is this sufficient evidence to reject future shipments from this supplier if the hospital’s quality standard requires 99.7 percent of the syringes to be acceptable?

(b)State the hypotheses and decision rule.
Ho: p = 0.997
Ha: p < 0.997 (evidence to reject supplier)
-------
test statistic:: z(0.995) = (0.995-0.997)/sqrt[0.997*0.003/2880] = -1.9625
Decision rule: Reject Ho is ts< -1.645
--------
(d) Find the p-value.
p-value = P(z<1.9625) = 0.0249
====================
Conclusion:
Since the p-value is less than 5%, reject Ho.
The test results support rejecting the supplier.
========
Cheers,
Stan H.
===========